OpenStudy (anonymous):

find an appropriate solution (to 4 decimal places) for each equation 3^x+2=5

OpenStudy (anonymous):

3^x = 5 -2 3^x = 3^1 ->x = 1

OpenStudy (anonymous):

wouldnt that make it if plug 1 in for x 3^3=5

OpenStudy (anonymous):

should be a number i have to round to 4 decimal places

OpenStudy (anonymous):

I probably a general instruction for all problems!

OpenStudy (anonymous):

what do you mean by that

OpenStudy (anonymous):

In this case, luckily the result is even, but most of other cases it's not!

OpenStudy (anonymous):

how can x=1

OpenStudy (anonymous):

put it back in original problem it comes out 3^3=5

OpenStudy (anonymous):

That's the whole entire solution already!

OpenStudy (anonymous):

that would be 27=5